Solve the system of equations.
1.5x – 7y = 2
0.625x + 3y = –2.125

Answer by Alan3354(69443)

Solve the system of equations.
1.5x – 7y = 2
0.625x + 3y = –2.125
----------
get rid of fractions for convenience
3x - 14y = 4 (x2)
5x + 24y = -17 (x8)
I'll use determinants
x y 1
3 -14 -4
5 24 17
Det = 3*24 - 5*(-14)
Det = 142
142x = -14*17 - 24*(-14)
= -238 + 96
= -142
x = -1
---------
-142y = 3*17 - 5*(-4)
= 51 + 20
=71
y = -71/142
y = -1/2